Matrix Methods/Linear Algebra: Please show all work and justify the answer!

8. Find the eigenvalues of each matrix. -4 2 (a) (8 points) A= 6 7 [ 1 (b) (4 points) A = 3 0 0 1 -2 0 2 3 4
Answer #1

8 (a) А [ - 2 - 4 20 6 7 The characteristic equation is - la-x. Il=0 → 115]-[:] 5) 11- => - 4- 6 2 7-1 1 -4-1 6 2 7-1 -(4+4)지 2831412=0 것 3 -40 = 0 키 3 I 9 -4-11-40) 2 1 = 25So the ergen values +1=8, 12=-5 (b) 3 O 2 3 4 The characteristic equation is 1 A Il=0 102487-1888] 3-doo -2-16 2 3. 4 -1 =) (